# Agentic Large Language Models, a survey

### v2

@@ 2.2 Self Reflection
Reasoning methods draw inspiration from step-by-step human solution approaches. The more elaborate approaches use explicit planning-like methods to look ahead and use feedback for verification. Such prompt-improvement loops facilitate a form of self reflection, since the model assesses and improves its own results. Self reflection happens when an external algorithm uses the LLM to assess its own predictions, and creates a new prompt for the same LLM to come up with a better answer. The improvement loop improves the prompts by using external memory, outside the LLM.333Note that self reflection generates new data that can be used for the model to train on. Whether the data is used for training depends on the training scheme: in-context learning does not update the model’s parameters, finetuning does. Note that in describing our taxonomy, we are now in the middle of the transition from passive model to active agent, as the agent is assessing its model’s predictions, and tries to improve them through reflection.
2.2.1 Prompt-Improvement
@@
Progressive hint prompting (PHP) is a reinforcement learning approach to interactively improve prompts (Zheng et al., 2023). PHP works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and a subsequent answer is obtained. We (3) repeat operation (2) until the answer becomes stable, just as the policy must converge in a regular policy-optimizing reinforcement learning algorithm. The authors have combined this approach with Chain of Thought and Self Consistency. Using GPT-4, state-of-the-art performance was achieved in grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%) (Zheng et al., 2023).
2.2.2 Using LLMs for Self Reflection
@@
The Self Refine approach is motivated by acquiring feedback from an LLM to iteratively improve the answers that are provided by that LLM (Madaan et al., 2023). In this approach, initial outputs from LLMs are used to improve the prompt through iterative feedback and refinement. Like PHP, the LLM generates an initial output and provides feedback for its answer, using it to refine itself, iteratively. Figure 7 illustrates the approach. Self-refine prompts the LLM in three ways: (0) for initial generation, (1) for feedback, and (2) for refinement. Figure 8 provides pseudo-code for the algorithm, in which the three calls to the LLM are clearly shown. The three prompts are labeled . (The equation numbers in the figure refer to the original paper.)
@@
Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked on dialogue response generation (Askari et al., 2024), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models.
@@
An earlier approach is ReAct (Yao et al., 2022), which has been further refined by (Shinn et al., 2024) as Reflexion. The goal is to create an agent that learns by reflecting on failures in order to enhance its results, much like humans do. Like Self Refine, Reflexion uses three language model prompts: an actor-LLM, an evaluator-LLM, and a reflector-LLM (which can be separate instances of the same model). Reflexion works as follows: (1) the actor generates text and actions, (2) the evaluator model scores the outputs produced by the actor, and (3) the self reflection model generates verbal reinforcement cues to assist the actor to self improve (see Figure 9).
@@
An approach called Self Discover goes a step further (Zhou et al., 2024a). This approach lets the agent analyze a problem, and discover which prompts work best. (It uses a dataset of prompts from a number of self reflective or chain of thought prompts, taken from PromptBreeder (Fernando et al., 2023).) The prompts are then adapted to the problem, and refined. Other approaches take a metalearning approach (Huisman et al., 2021). Buffer of Thoughts (Yang et al., 2024c) and Meta Chain of Thought (Xiang et al., 2025) extend traditional Chain of Thought by explicitly modeling the underlying reasoning required to arrive at a particular chain of thought. Further self reflection approaches that are based on external reflection algorithms are reviewed by Plaat et al. (2024).
Intrinsic Reasoning
In contrast to explicit reasoning algorithms that are external to the model, intrinsic reasoning is performed by the model itself, the model has integrated reasoning capabilities in its (trained) architecture rather than relying on external prompts and methods. In Self taught reasoner (Zelikman et al., 2022) data is generated by reasoning at inference time, that is then used to augment supervised finetuning training data.
A related approach was proposed in the development of DeepSeek-R1 Guo et al. (2025). This method distinguishes itself from external reasoning approaches by emphasizing the model’s self-generated reasoning steps. It learned the steps through reinforcement learning, integrating data generation and training in one loop. This intrinsic approach holds significant potential for creating more autonomous and adaptive AI systems. Lowering the supervised data requirements to train LLMs, DeepSeek’s methodology leverages reinforcement learning to enable models like DeepSeek-R1-Zero (the reasoning LLM before preference fine-tuning) to evolve reasoning skills autonomously. By starting with a base model and applying reinforcement learning, the system identifies and reinforces effective reasoning patterns. DeepSeek uses GRPO, which eliminates the need for a separate critic model by calculating advantages with group-based scoring (Shao et al., 2024). This process allows the model to explore various problem-solving strategies and refine its thought processes without external inference-time control loops.
One of the key features of DeepSeek’s approach is the emergence of sophisticated reasoning behaviors, such as reflection and exploration of alternative problem-solving methods (Mercer et al., 2025). These behaviors arise spontaneously as a result of the model’s interaction with the reinforcement learning environment, rather than being pre-programmed. For example, DeepSeek-R1-Zero learns to allocate more thinking time to problems by reevaluating its initial approach. This autonomous approach of learning to reason could lead to more stable and adaptive reasoning LLMs.
@@
Note that the transformer architecture has been proven to be able to simulate Turing machines (Pérez et al., 2021), and therefore, in theory, the prompt-improvement loop, and the memory, could be implemented inside the transformer itself, internal to the LLM. Some studies pursue this idea further, and see how the current external control algorithms can be made internal. Schultz et al. (2024) showed how LLMs can be trained to do a tree search. Giannou et al. (2023) show how programs written in Restricted Access Sequence Processing language (RASP) can be mapped onto transformer networks. They show how looped transformers (transformers whose input neurons are connected to their output neurons) can emulate a basic calculator, a basic linear algebra library, and in-context learning algorithms that employ back-propagation. This is an area for further research.
@@
In general, the use of memory between prompts allows individual LLMs to acquire experience. The prompt history of a model determines individual preferences, anthropomorphically speaking, the agent acquires a personality. Zhang et al. (2023b) studies how memory coordination is an element for LLM personalization. In another study, Think in Memory is an architecture to model human-like memory processes to selectively recall historical thoughts in long term interaction scenarios (Liu et al., 2023).
2.3 Retrieval Augmentation
@@
It is now time to review works that address another shortcoming of LLMs, the lack of timely information. Retrieval augmentation improves models by including information of timely or specialized nature, that was not yet available during pretraining. This can be stock data, a recent hotel booking, or data that has to be retrieved from specialized databases that was not included in the pretraining corpus. Retrieval of such data is usually done at inference time, with tools from the field of databases (Cong et al., 2024), information retrieval (Verberne, 2010; Baeza-Yates et al., 1999), and knowledge representation (Van Harmelen et al., 2008).
An LLM is pretrained on a large text corpus. The LLM’s parameters can not contain information that was generated after the corpus was fixed. Inference-time lookup must be added if the LLM is to answer timely or specialized queries. How can such information be added to an LLM?
@@
Retrieval augmented generation methods (RAG) have been developed to integrate data from external knowledge bases with the pretrained knowledge stored in the parameters of language models. Most RAG approaches work on unstructured (textual) data sources. These text documents are indexed to increase efficient access, and they can be organized and a knowledge graph can be created. Finally, database-type query optimization is often performed, where queries can be expanded or complex queries can be split into sub-queries (Cong et al., 2024).
@@
Adaptive retieval methods enable LLMs to determine the optimal moment for retrieval (Asai et al., 2023). These methods are related to self reflection (see Section 2.2), so that LLMs employ inference-time judgment on when and how to retrieve. For example, Graph Toolformer (Zhang, 2023) applies techniques from Self Ask (Press et al., 2022) to initiate search queries, allowing the LLM to decide when to retrieve extra information.
The approach by Lewis et al. (2020) augments pre-trained LLMs with knowledge from different knowledge bases. This information is stored in a dense vector index of Wikipedia. Both components are finetuned in a probabilistic model that is trained end-to-end.
@@
The integration of time sensitive unstructured (information retrieval) and structured (database) data with LLMs is a fruitful and important area for agentic LLM. Gao et al. (2023c) review many different RAG approaches.
2.4 Discussion
@@
In this section, we have surveyed techniques that have been developed to improve decision making by LLMs. We have seen that modern approaches are neurosymbolic: the deep learning AI tradition (neural networks and transformers) is joined at inference-time by the symbolic AI tradition (reasoning, planning and knowledge retrieval). We discussed the need for better reasoning performance, which started with solving math word problems. We discussed how research on reasoning LLMs was jump-started by the Chain of Thought paper (Wei et al., 2022b), where a single addition to the prompt caused the model to perform implicit step-by-step reasoning. Inference-time planning and search algorithms perform explicit prompt improvement, using memory to keep track of the current prompt (Ko et al., 2024; Giannou et al., 2023). Self-prompt-improvement approaches perform self reflection; the memory of individual prompt histories gives models their own history, and, perhaps, personality. Self reflection in LLMs is related to theory of mind, see Section 4.1.3. Self reflection methods have achieved much attention and success.
Furthermore, in some use cases LLMs need up to date information, beyond that which was available in their training corpus (Miikkulainen, 2024). Retrieval augmented generation is an active field that accesses specialty knowledge bases and search engines (such as Google or Wikipedia). Indexing and knowledge graphs are some of the knowledge retrieval techniques that are used (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2023c). RAG uses tools to access information, at inference time, outside the LLM. RAG methods also use self reflection methods to determine the best moment for retrieval (Asai et al., 2023; Press et al., 2022). Further surveys are Shen (2024); Li (2024). The use of tools creates a bridge to the next category of the taxonomy: LLMs that act in the outside world.
Much research has been performed on decision making and reasoning by LLMs. New data is generated by retrieval, and by the use of tools. However, prompt learning methods do not change the parameters of the model; in order to use the data that is generated by inference-time approaches, finetuning must be used.
2.4.1 Thinking, Fast and Slow
@@
In 2011 Kahneman published the book Thinking, Fast and Slow in which the terms System 1 and System 2 were used to distinghuish human thought into intuitive, fast, thinking, and deliberative, slow, thinking (Kahneman, 2011). These terms have become popular in artificial intelligence. At inference time, pure LLMs think fast (System 1 thinking). Inference-time step-by-step methods can be added to achieve deliberative slow thinking (System 2 thinking). LLMs are based on the deep learning AI-tradition (System 1 thinking). The use of tools at inference time enhances the LLM part with knowledge retrieval or processing tools from the symbolic AI tradition (System 2 thinking).
@@
We should note that whereas researchers sometimes humanize LLMs and their capabilities, LLMs only perform next-token prediction. By generating more tokens to form an answer (reasoning step-by-step), the token-path from the prompt to the final answer becomes longer. The reason that this leads more often to correct answers, might be because it takes smaller steps into the direction of the answer, making the correct answer more plausible with every step in between. Reasoning is narrowing down probabilities such that the correct answer becomes more probable to generate, independent of interpretations related to human cognition (Guo et al., 2025).
2.4.2 Artificial General Intelligence

### v3

@@ 2.2. Self Reflection
Self reflection happens when an external algorithm uses the LLM to assess its own predictions, and creates a new prompt for the same LLM to come up with a better answer (see for example the algorithm in Figure 9). The improvement loop improves the prompts by using external memory, outside the LLM.333Note that self reflection generates new data that can be used for the model to train on. Whether the data is used for training depends on the training scheme: in-context learning does not update the model’s parameters, finetuning does. Note that in describing our taxonomy, we are now in the middle of the transition from passive model to active agent, as the agent is assessing its model’s predictions, and tries to improve them through reflection.
2.2.1. Prompt-Improvement
@@
Progressive hint prompting (PHP) is a reinforcement learning approach to interactively improve prompts (zheng2023progressive). PHP works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and a subsequent answer is obtained. We repeat operation (2) until the answer becomes stable, just as the policy must converge in a regular policy-optimizing reinforcement learning algorithm. The authors have combined this approach with Chain of Thought and Self Consistency. Using GPT-4, state-of-the-art performance was achieved in grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%) (zheng2023progressive).
2.2.2. Using LLMs for Self Reflection
@@
The Self Refine approach is motivated by acquiring feedback from an LLM to iteratively improve the answers that are provided by that LLM (madaan2023self). In this approach, initial outputs from LLMs are used to improve the prompt through iterative feedback and refinement. Like PHP, the LLM generates an initial output and provides feedback for its answer, using it to refine itself, iteratively. Figure 8 illustrates the approach. Self-refine prompts the LLM in three ways: (0) for initial generation, (1) for feedback, and (2) for refinement. Figure 9 provides pseudo-code for the algorithm, in which the three calls to the LLM are clearly shown. The three prompts are labeled . (The equation numbers in the figure refer to the original paper.)
@@
Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked on dialogue response generation (askari2024self), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models.
@@
An earlier approach is ReAct (yao2022react), which has been further refined by (shinn2024reflexion) as Reflexion. The goal is to create an agent that learns by reflecting on failures in order to enhance its results, much like humans do. Like Self Refine, Reflexion uses three language model prompts: an actor-LLM, an evaluator-LLM, and a reflector-LLM (which can be separate instances of the same model). Reflexion works as follows: (1) the actor generates text and actions, (2) the evaluator model scores the outputs produced by the actor, and (3) the self reflection model generates verbal reinforcement cues to assist the actor to self improve (see Figure 10).
@@
An approach called Self Discover goes a step further (zhou2024self). This approach lets the agent analyze a problem, and discover which prompts work best. (It uses a dataset of prompts from a number of self reflective or chain of thought prompts, taken from PromptBreeder (fernando2023promptbreeder).) The prompts are then adapted to the problem, and refined. Other approaches take a metalearning approach (huisman2021survey). Buffer of Thoughts (yang2024buffer) and Meta Chain of Thought (xiang2025towards) extend traditional Chain of Thought by explicitly modeling the underlying reasoning required to arrive at a particular chain of thought. Further self reflection approaches that are based on external reflection algorithms are reviewed by plaat2024reasoning.
@@
Note that the transformer architecture has been proven to be able to simulate Turing machines (perez2021attention), and therefore, in theory, the prompt-improvement loop, and the memory, could be implemented inside the transformer itself, internal to the LLM. Some studies pursue this idea further, and see how the current external control algorithms can be made internal. schultz2024mastering showed how LLMs can be trained to do a tree search. giannou2023looped show how programs written in Restricted Access Sequence Processing language (RASP) can be mapped onto transformer networks. They show how looped transformers (transformers whose input neurons are connected to their output neurons) can emulate a basic calculator, a basic linear algebra library, and in-context learning algorithms that employ back-propagation. This is an area for further research (li2025implicit).
@@
In general, the use of memory between prompts allows individual LLMs to acquire experience. The prompt history of a model determines individual preferences, or, anthropomorphically speaking, the agent acquires a personality. zhang2023memory study how memory coordination is an element for LLM personalization. In another study, Think in Memory, is an architecture to model human-like memory processes to selectively recall historical thoughts in long term interaction scenarios (liu2023think).
Implicit Reasoning
In contrast to explicit reasoning algorithms that are external to the model, implicit reasoning is performed by the model itself, the model has integrated reasoning capabilities in its (trained) architecture rather than relying on external prompts and methods. In Self taught reasoner (zelikman2022star) data is generated by reasoning at inference time, that is then used to augment supervised finetuning training data. The field of implicit reasoning is an active area of research. For a survey, see li2025implicit.
A related approach was proposed in the development of DeepSeek-R1 (guo2025deepseek). This method distinguishes itself from external reasoning approaches by emphasizing the model’s self-generated reasoning steps. It learned the steps through reinforcement learning, integrating data generation and training in one loop. This intrinsic approach holds significant potential for creating more autonomous and adaptive AI systems. Lowering the supervised data requirements to train LLMs, DeepSeek’s methodology leverages reinforcement learning to enable models like DeepSeek-R1-Zero (the reasoning LLM before preference fine-tuning) to evolve reasoning skills autonomously. By starting with a base model and applying reinforcement learning, the system identifies and reinforces effective reasoning patterns. DeepSeek uses group relative policy optimization (GRPO), which eliminates the need for a separate critic model by calculating advantages with group-based scoring (shao2024deepseekmath). This process allows the model to explore various problem-solving strategies and refine its thought processes without external inference-time control loops. A popular related method is reinforcement learning with verifiable rewards (RLVR), which also uses rewards that can be quickly calculated for chains of thought for finetuning the model (lambert2024tulu).
One of the key features of these approaches is the emergence of sophisticated reasoning behaviors, such as reflection and exploration of alternative problem-solving methods (mercer2025brief). These behaviors arise spontaneously as a result of the model’s interaction with the reinforcement learning environment, rather than being pre-programmed. For example, DeepSeek-R1-Zero learns to allocate more thinking time to problems by reevaluating its initial approach. This autonomous approach of learning to reason could lead to more stable and adaptive reasoning LLMs.
2.3. Retrieval Augmentation
@@
Another shortcoming of LLMs is the lack of timely information. Retrieval augmentation improves models by including information of a timely or specialized nature, that was not yet available during pretraining. This can be stock data, a recent hotel booking, or data that has to be retrieved from specialized databases that was not included in the training corpus. Retrieval of such data is usually done at inference time, with tools from the field of databases (cong2024query), information retrieval (verberne2010search; baeza1999modern), and knowledge representation (van2008handbook).
@@
Most retrieval augmented generation methods (RAG) work on unstructured (textual) data sources. These text documents are indexed to increase efficient access, and can be organized as a knowledge graph. Furthermore, database-type query optimization is often performed, where queries can be expanded or complex queries can be split into sub-queries (cong2024query).
@@
Adaptive retrieval methods enable LLMs to determine the optimal moment for retrieval (asai2023self). These methods are related to self reflection (see Section 2.2). For example, Graph Toolformer (zhang2023graph) applies techniques from Self Ask (press2022measuring) to initiate search queries, allowing the LLM to decide when to retrieve extra information. The approach by lewis2020retrieval augments pre-trained LLMs with information from different knowledge bases, such as Wikipedia. This information is stored in a dense vector index. Both components are finetuned in a probabilistic model that is trained end-to-end.
@@
Retrieval augmentation can be costly. Researchers are looking into combining curated ground truth with synthetic data, with the LLM in the role of judge and self evaluator (vanelburg2025evaluateragssyntheticdata; es-etal-2024-ragas). The integration of time sensitive unstructured (information retrieval) and structured (database) data with LLMs is a fruitful and important area for agentic LLM. gao2023retrieval review many different RAG approaches. Further surveys are shen2024llm; li2024review.
2.4. Discussion
@@
In this section, we have surveyed techniques that have been developed to improve decision making by LLMs. We discussed the need for better reasoning performance, which started with solving math word problems. Modern approaches are neurosymbolic: the deep learning AI tradition (neural networks and transformers) is joined at inference-time by the symbolic AI tradition (reasoning, planning and knowledge retrieval).
We have surveyed individual reasoning methods. In order to dig deeper into the reasoning foundations of agentic LLM, we will now discuss Chain of Thought and Self Reflection in more detail.
2.4.1. In Depth: Chain of Thought and Self Reflection
Chain of Thought
Research on multi-step reasoning LLMs was jump-started by the Chain of Thought paper (wei2022chain), where a single addition to the prompt caused the model to perform implicit step-by-step reasoning. Chain of Thought has lead to a strong increase of in-context reasoning performance by LLMs. For agents that interact in the real world it is important to be able to perform multi-step reasoning tasks and to interact with other agents. Chain of Thought has been instrumental in this respect.
To augment finetuning for math reasoning and coding tasks, reinforcement learning with verifiable rewards, RLVR, (lambert2024tulu) and group relative policy optimization, GRPO, (guo2025deepseek) use inference-time Chain of Thought reasoning traces. These methods are reportedly used by OpenAI o1 and o3 (wu2024comparative), DeepSeek (guo2025deepseek), and Qwen-3 (yang2025qwen3). RLVR and GRPO have created the possibility to trade off training time to model size, allowing resource-efficient training of LLMs. muennighoff2025s1 analyze how such test-time scaling can trade-off model size and training time.
Self Reflection
Agents that interact should reflect on their own behavior, in order to adapt and learn. Planning allows an agent to imagine possible futures, allowing LLMs to interact in a more intelligent way in the real world. We have discussed various inference-time self reflection methods that used planning and search algorithms and perform explicit prompt improvement (ko2024memreasoner; giannou2023looped).
Self reflection in LLMs is related to theory of mind. Self reflective methods allow LLMs to reason about expected behavior of the agents that it interacts with (see also Section 4.1.3).
We should note that self reflection is not without its challenges. In self reflective approaches the same LLMs are used in two or more ways, for example to generate subproblems, and to evaluate them. When errors arise, there are now two (or more) types of prompts to test. Furthermore, even when the individual prompts work as expected, they may interact in unexpected ways. Debugging self reflective agents can be challenging. wang2025hierarchical introduce hierarchical reasoning models, where the different models learn at different speeds, in an attempt to reduce oscillations between the two interacting learning models. A second problem with self reflection is that after many interactions, the traces of states, actions and rewards can become too long for the context window (liu277271533comprehensive). Various long-context models and methods to compress the traces have been proposed (li2025flow; zhang2025agentic).
Self reflection is a promising, but challenging, area of research, that is of great importance to agentic LLMs.
2.4.2. Thinking, Fast and Slow
@@
In 2011 Kahneman published the book Thinking, Fast and Slow in which the terms System 1 and System 2 were used to distinguish human thought into intuitive, fast, thinking, and deliberative, slow, thinking (kahneman2011thinking). These terms have become popular in artificial intelligence. At inference time, pure LLMs think fast (System 1 thinking). Inference-time step-by-step methods can be added to achieve deliberative slow thinking (System 2 thinking). LLMs are based on the deep learning AI-tradition (System 1 thinking). The use of tools at inference time enhances the LLM part with knowledge retrieval or processing tools from the symbolic AI tradition (System 2 thinking).
@@
We should note that whereas researchers sometimes humanize LLMs and their capabilities, LLMs only perform next-token prediction. By generating more tokens to form an answer (reasoning step-by-step), the token-path from the prompt to the final answer becomes longer. The reason that this leads more often to correct answers, might be because it takes smaller steps into the direction of the answer, making the correct answer more plausible with every step in between. Reasoning is narrowing down probabilities such that the correct answer becomes more probable to generate, independent of interpretations related to human cognition (guo2025deepseek; lambert2023measuring).
2.4.3. Causal and Common Sense Reasoning
While LLMs exhibit logical reasoning, a key limitation lies in the domain of deep comprehension. For instance, as a play on stochastic parrots (bender2021stochastic), LLMs are often criticized as being causal parrots that are good at reproducing causal language from their training data but lack true causal inference capabilities. LLMs struggle with abstract or counterfactual reasoning, necessary for robust decision-making (zevcevic2023causal; chi2024unveiling). Furthermore, the use of the terms reasoning and thinking has been questioned, in a study highlighting that current reasoning approaches fail to solve modestly complex puzzle problems such as Towers of Hanoi (shojaee2025illusion). Although the study has been criticized (opus2025illusion), the outcome that LLMs do not perform well on combinatorial puzzles has been replicated (paglieri2024balrog; ruoss2024lmact; su2025limits).
Similarly, for agents to act appropriately, a degree of common sense reasoning is required. This also remains a challenge, as LLMs frequently struggle when tested on abstract common sense tasks (zhou2020evaluating). Overcoming these gaps is crucial for developing agents that are reliable in real-world environments, for example, to prevent that simple adverserial injection of irrelevant factoids in a prompt can cause a reasoning model to overthink a problem by up to 50% and substantially impact error rates (rajeev2025catsconfusereasoningllm). Especially in an agentic setting this can be problematic, as much of the context is generated, leading to longer and weaker contexts.
2.4.4. Artificial General Intelligence